# Erased, But Not Forgotten: Erased Rectified Flow Transformers Still Remain Unsafe Under Concept Attack

### v2

@@ 6 Conclusion
In this work, we introduce ReFlux, the first systematic concept attack method for rectified flow transformers. Our analysis explains why certain prompt-based strategies effective in SD fail to transfer, uncovering that erasure fundamentally operates through attention localization. Building on this insight, we propose a lightweight LoRA-based fine-tuning approach with stabilized reverse-attention optimization, velocity-guided dynamics, and consistency constraints, enabling precise reactivation of erased concepts while preserving global layout and unrelated content. Extensive experiments demonstrate not only stronger attack performance but also high-fidelity generation and reliable preservation of irrelevant elements. Beyond surpassing baselines, ReFlux establishes the first reliable and extensible benchmark for evaluating the robustness of concept erasure in next-generation flow matching T2I framework.
Ethics Statement
Our work introduces ReFlux, an attack method explicitly designed to evaluate the robustness of safety-oriented concept erasure in next-generation rectified flow T2I models. Our goal is to provide a systematic evaluation of erasure defenses, not to promote harmful content generation. All experiments are conducted offline on public available datasets and detection classifiers, with potentially sensitive outputs blurred or down-scaled.
We acknowledge the dual-use nature of this research: while the attack could, in principle, regenerate restricted concepts, we adopt multiple concrete mitigations. We will release the full training code and detector weights publicly to ensure reproducibility, but will not include any pretrained models that directly reproduce high-risk content. Released artifacts will default to safe configurations: built-in content filters enabled, integration examples that call platform moderation APIs (e.g., Huggingface safety checkers), and sanitized demo assets. The repository will include a clear responsible-use licence and an explicit README describing safe defaults, human-in-the-loop recommendations, and procedures for requesting controlled access to any sensitive materials for vetted researchers. We also recommend runtime mitigations (rate limiting, API gating, and content-moderation pipelines) for anyone deploying these tools.
By revealing that many current erasure strategies on Flux often suppress concepts only superficially, our findings aim to inform the development of stronger defenses, safer deployment practices, and more reliable evaluation standards for generative models. We commit to the ICLR code of ethics and call on the community to advance dataset purification, safety alignment, and interpretability research to ensure the responsible progress of generative AI.
Reproducibility statement
We have taken extensive measures to ensure the reproducibility of our work. Section 5.1 outlines the experimental setup, training objectives, hardware environment, and baseline configurations, with additional implementation details and adopted hyperparameters provided in Appendix F. All datasets used in our study, including I2P, ConceptPrune, and CelebA, are publicly accessible, and the full lists of additional benchmarks (e.g., Relationship, Abstraction) are provided in Appendix H. To further facilitate reproducibility, we include the complete source code, all experiment scripts, and classifier weights (with no unsafe content) in the Supplementary materials, and we will make the full codebase and resources publicly available upon publication.
@@ Appendix A Rectified Flow Transformers and Flux Architecture
Rectified flow transformers are a new class of T2I models that replace the iterative denoising process of diffusion with a flow matching objective that directly learns the trajectory between data and noise distributions (Lipman et al., 2022). Instead of relying on stochastic score estimation and noise schedules, rectified flows provide a deterministic path, which improves convergence stability, reduces sampling complexity, and yields stronger alignment with conditioning signals. Architecturally, these models depart from the U-Net backbone of SD and adopt a transformer-based design that better integrates text and image information, echoing recent advances in large-scale language modeling. This shift not only leads to improved sample quality but also introduces new structural mechanisms that are central to our analysis of concept erasure and attack.
@@
Figure 7: Architecture of Flux [dev]. Flux [dev] employs frozen CLIP-L/14(Radford et al., 2021) and T5-XXL(Raffel et al., 2020) as text encoders for caption feature extraction. The coarse CLIP embedding, concatenated with the timestep embedding $y$, enters the modulation pathway, while the fine-grained T5 embedding $c$ is concatenated with the noised image latents $x$. These fused representations are processed through nineteen dual stream blocks and thirty-eight single stream blocks to predict outputs in the VAE latent space. Within each dual stream block, concatenation of projections implicitly replaces explicit cross-attention, allowing token-level semantics to emerge as localized heatmaps. This mechanism provides the structural basis for concept erasure and reactivation in our study.
Following this discovery, we direct our optimization efforts within the dual stream block. Careful experimentation shows that Add_v_proj and To_v are numerically unstable and thus unsuitable for controlled optimization. In contrast, query and key pathways (Add_q_proj, Add_k_proj, To_q, To_k) prove both stable and sensitive, offering reliable entry points for manipulating semantic alignment. This insight underpins our method design, ensuring that interventions suppress or reactivate concepts without destabilizing global generation. For fairness, we adapt existing baselines originally designed to optimize Q and V in SD (e.g., ESD (Gandikota et al., 2023), AC (Kumari et al., 2023))—to instead optimize Q and K within Flux. This adjustment provides a consistent comparison framework and demonstrates how architectural differences in rectified flow transformers necessitate targeted methodological changes.
@@ Appendix B Difference Between T5 and CLIP Embeddings
A fundamental difference between the text encoders in Flux and SD lies in their training paradigms and representational granularity. T5 is trained on large-scale textual corpora with a sequence-to-sequence objective, emphasizing sentence-level semantics and contextual dependencies (Raffel et al., 2020). Its embeddings are optimized for holistic sentence meaning rather than fine-grained concept separation, which makes them sensitive to contextual co-occurrence in natural language. By contrast, CLIP is trained with a multimodal contrastive learning objective (Radford et al., 2021), where text and image pairs are aligned and mismatches are separated. This training paradigm structures the semantic space according to visually grounded concepts, leading to embeddings that reflect a different notion of similarity. These distinctions explain why Flux, relying on T5 embeddings, exhibits different behavior from SD.
@@ Appendix C Attention Localization under Erasure Methods
Figure 8: Visualization of attention localization in concept erasure. The attention heatmaps confirm that existing erasure strategies (Gao et al., 2025; Kumari et al., 2023; Gandikota et al., 2023; Bui et al., 2024) share a common mechanism: attention localization. Localized regions corresponding to target concepts are suppressed, visually corroborating our theoretical analysis.
@@ Appendix D Why Naive Attention Maximization Diverges
$p_{i}=\frac{e^{z_{i}}}{\sum_{j=1}^{m}e^{z_{j}}},\quad i=1,\dots,m,\qquad\text{so }p\in(0,1)^{m},\ \sum_{i}p_{i}=1.$ (9)
@@
$z_{k}-\max_{j\neq k}z_{j}\to+\infty\quad\Longrightarrow\quad p_{k}\to 1,\ p_{j\neq k}\to 0.$ (11)
@@
is structurally prone to divergence and vanishing gradients—precisely the recipe for the “attention divergence” we observe in Figure 2 (c).
@@ Appendix E Flux Can Generate Consistent Content from Shuffled Prompts
The issue of overfitting arises from fixing the token_id of the target concept. To mitigate this, we introduce dynamic variation of the target token_id across training iterations. Prior work Gao et al. (2025) speculated that randomly shuffling the prompt should not affect the generation quality of Flux. We extend this hypothesis with concrete evidence.
@@ Appendix F Baselines and Implementation Details
Erase AC (Kumari et al., 2023)
ESD (Gandikota et al., 2023)
EraseAnything (Gao et al., 2025)
EAP (Bui et al., 2024)
ConceptPrune (Chavhan et al., 2025)
Attack P4D (Chin et al., 2024)
UnlearnDiffAtk (Zhang et al., 2024)
Ring-A-Bell (Tsai et al., 2024)
Reason2Attack (Zhang et al., 2025)
@@ F.1 Baselines for Attack Evaluation
We adopt UnlearnDiffAtk (Zhang et al., 2024) as a representative white-box attack baseline, which leverages gradient-based noise prediction loss to optimize adversarial prompts, outperforming prior methods such as P4D (Chin et al., 2024). For the black-box setting, we include Ring-A-Bell (Tsai et al., 2024) and its enhanced variant Ring-A-Bell-Union, both of which construct a target concept direction from positive–negative prompt pairs and inject it into the prompt embedding to revive the forgotten concept, demonstrating stronger performance compared with approaches like QF-Attack (Zhuang et al., 2023). For emerging methods, we consider the emerging LLM-based method Reason2Attack Zhang et al. (2025) as a representative of reasoning-driven attack strategies.
@@ F.2 Concept Erasure Methods for Evaluation
We select publicly accessible and reproducible concept erasure methods as victim models for evaluation. This includes classical approaches like ESD (Gandikota et al., 2023) and AC (Kumari et al., 2023), as well as more recent methods, including EAP (Bui et al., 2024) and EraseAnything (EA) (Gao et al., 2025), the latter being the first approach tailored for rectified flow transformers. For ESD under nudity and violence settings, we fine-tune both non-cross-attention and cross-attention parameters with negative guidance factors of 1 and 3, respectively. We exclude UCE (Gandikota et al., 2024), as its overly aggressive removal severely distorts Flux outputs. All baselines and ablated models follow official implementations.
@@ H.2 Complete list of Entity, Abstraction, Relationship
This dataset is augmented on Gao et al. (2025), covering more abstract and diverse test categories. The full list used in our experiments is presented in Table 8.
@@ H.3 Implementation Details of the Celebrity Benchmark
To construct a reliable benchmark for evaluating celebrity-related erasure and attack methods, we curate a refined subset from the CelebA dataset (Liu et al., 2018). During this process, we deliberately exclude those individuals that Flux [dev] is unable to faithfully reconstruct. A manual inspection procedure is applied, where we compare synthesized outputs against their textual prompts and further supplement the pool with several well-known comic characters. This selection process ultimately yields a dataset of 100 celebrities, which we evenly divide into two groups: 50 designated for attack evaluation and 50 retained as control cases. The specific names of the celebrities used in our ablation study are listed in Table LABEL:table:_appendix_celebrity.
@@
For recognition tasks, we implement a lightweight yet effective classification network based on MobileNetV2 (Howard et al., 2017) pretrained on ImageNet (Deng et al., 2009). On top of the original architecture, we append a GlobalAveragePooling2D layer followed by a fully connected Softmax layer. Training is performed with the Adam optimizer using a fixed learning rate of 1e-4, and categorical cross-entropy is adopted as the loss function. For training dataset, we first generate 50 images per celebrity (fixed prompt and random seeds), amounting to a total of 5,000 images. We then randomly re-sample the dataset and partitioned it into training (80%) and testing (20%) splits.
@@ H.4 Additional Results
Benchmarking Against State-of-the-Art (SOTA). Figure 11 shows the result of benchmarking ReFlux against representative SOTAs erasure attacks on the I2P dataset. For evaluation, we adopt NudeNet with a detection threshold of 0.6, which is commonly regarded as a reasonable boundary for identifying sensitive content. It should be noted, however, that surpassing this threshold does not necessarily imply the actual exposure of body organs, but rather indicates that the generated image triggers the detector’s nudity confidence. Among baselines, UnlearnDiffAtk exhibits a notable failure mode when applied to rectified flow models: it frequently collapses into low-quality or distorted generations. Nevertheless, due to the coarse sensitivity of the NudeNet classifier, such degraded outputs are still often flagged as “nude”, inflating its measured attack success. By contrast, our approach not only achieves higher quantitative scores under the same detector but also maintains high-fidelity and semantically consistent generations, providing a more accurate and reliable benchmark of erasure robustness.
@@
Figure 12: More visualization on artistic style (“Pablo Picasso”) attacks from the Chavhan et al. (2025) dataset. While AC (Kumari et al., 2023) erasure removes the distinctive Picasso style, our method successfully restores the erased artistic patterns across diverse works.
Attacking Artistic Style Concepts. Artistic styles are representative benchmarks of abstract concepts, widely evaluated in concept erasure and attack. Here, we present more visualization results. Figure 12 presents results for Picasso-style prompts. The top row shows baseline generations from Flux.1 [dev], which faithfully capture Picasso’s unique color palette, geometric distortions, and expressive brushwork. After applying AC (Kumari et al., 2023) erasure (middle row), these stylistic signatures vanish almost entirely, producing outputs that resemble conventional photographic or illustrative imagery rather than Picasso’s style. Our attack (bottom row) effectively restores the erased style, reintroducing signature characteristics such as fragmented spatial composition, bold outlines, and vibrant color schemes. Importantly, the restored images not only achieve a high attack success rate but also preserve layout and semantic content from the erased generations (e.g., maintaining the same subject matter and scene structure). These results demonstrate that our method can reactivate highly abstract and global concepts, highlighting the robustness and generality of our approach. This also reveals that defenses like AC achieve only surface-level erasure, leaving deep conceptual residues that can be readily reawakened—“erased, but not forgotten”.
@@ H.5 User Study and VLM-based Evaluation
Human User Study. To obtain reliable human judgments, we design questionnaires that sampled from 7 evaluation categories: nudity, violence, artistic style, entity, abstraction, relationship, and celebrity. For each category, we randomly selected 5 sets of comparison results in our main experiments, yielding one complete questionnaire. Sensitive content such as nudity or violence is masked with bars or blur to ensure participant safety. In total, we recruited 20 non-artist participants, each of whom completed on average 10 questionnaires.
VLM-based Evaluation. We leverage the latest GPT-5222https://openai.com/index/introducing-gpt-5/ VLM as an automated evaluator, chosen for its strong reasoning ability, nuanced understanding of semantics, and robust visual grounding. For each of the 7 categories, we select 10 representative comparison sets and present them to GPT-5. The evaluation criteria are aligned with those used in the human study and specified in Table 10, which we format as structured user prompts.
As shown in Figure 16 and Figure 17, both our human user study and the GPT-5 VLM-based evaluation reveal a consistent trend. Our method outperforms all baseline approaches across the four metrics. Although Ring-A-Bell occasionally reports slightly higher values on irrelevant preservation, this is primarily because it fails to reactivate the target concept, producing outputs that remain almost indistinguishable from the erased baseline. In contrast, traditional PGD-based methods such as UnlearnDiffAtk and P4D often suffer from poor generation quality on Flux, exhibiting typical diffusion artifacts such as grid-like patterns, structural collapse, and mode instability, which result in substantially lower image quality scores. By effectively restoring the target concept while preserving prompt fidelity and visual coherence, our approach achieves both stronger semantic reactivation and more stable generative behavior, demonstrating clear advantages in robustness and reliability.
Figure 16: Human evaluation results across different attack methods. Bars show mean scores on a 1–5 scale, with error bars indicating standard deviations.

### v3

@@ 6 Conclusion
In this work, we introduce ReFlux, the first systematic concept attack method for rectified-flow transformers. We explain why prompt-based strategies that work for SD fail to transfer, showing that Flux erasure is driven by localized attention suppression. Building on this insight, we introduce a lightweight LoRA-based tuning strategy with stabilized reverse-attention optimization, velocity-guided dynamics, and consistency constraints, enabling precise concept reactivation while keeping layout and unrelated content intact. Experiments demonstrate strong attack performance and high-fidelity preservation. Beyond surpassing baselines, ReFlux establishes the first reliable and extensible benchmark for evaluating the robustness of concept erasure in next-generation flow matching T2I framework.
Supplementary Material
@@ Appendix A Rectified Flow Transformers and Flux Architecture
Rectified flow transformers are a new class of T2I models that replace the iterative denoising process of diffusion with a flow matching objective that directly learns the trajectory between data and noise distributions [24]. Instead of relying on stochastic score estimation and noise schedules, rectified flows provide a deterministic path, which improves convergence stability, reduces sampling complexity, and yields stronger alignment with conditioning signals. Architecturally, these models depart from the U-Net backbone of SD and adopt a transformer-based design that better integrates text and image information, echoing recent advances in large-scale language modeling. This shift not only leads to improved sample quality but also introduces new structural mechanisms that are central to our analysis of concept erasure and attack.
@@
Figure 8: Architecture of Flux ([dev] and [schnell] share the same architecture). Flux [dev] employs frozen CLIP-L/14[35] and T5-XXL[36] as text encoders for caption feature extraction. The coarse CLIP embedding, concatenated with the timestep embedding $y$, enters the modulation pathway, while the fine-grained T5 embedding $c$ is concatenated with the noised image latents $x$. These fused representations are processed through nineteen dual stream blocks and thirty-eight single stream blocks to predict outputs in the VAE latent space. Within each dual stream block, concatenation of projections implicitly replaces explicit cross-attention, allowing token-level semantics to emerge as localized heatmaps. This mechanism provides the structural basis for concept erasure and reactivation in our study.
Following this discovery, we direct our optimization efforts within the dual stream block. Careful experimentation shows that Add_v_proj and To_v are numerically unstable and thus unsuitable for controlled optimization. In contrast, query and key pathways (Add_q_proj, Add_k_proj, To_q, To_k) prove both stable and sensitive, offering reliable entry points for manipulating semantic alignment. This insight underpins our method design, ensuring that interventions suppress or reactivate concepts without destabilizing global generation. For fairness, we adapt existing baselines originally designed to optimize Q and V in SD (e.g., ESD [12], AC [20])—to instead optimize Q and K within Flux. This adjustment provides a consistent comparison framework and demonstrates how architectural differences in rectified flow transformers necessitate targeted methodological changes.
@@ Appendix B Difference Between T5 and CLIP Embeddings
A fundamental difference between the text encoders in Flux and SD lies in their training paradigms and representational granularity. T5 is trained on large-scale textual corpora with a sequence-to-sequence objective, emphasizing sentence-level semantics and contextual dependencies [36]. Its embeddings are optimized for holistic sentence meaning rather than fine-grained concept separation, which makes them sensitive to contextual co-occurrence in natural language. By contrast, CLIP is trained with a multimodal contrastive learning objective [35], where text and image pairs are aligned and mismatches are separated. This training paradigm structures the semantic space according to visually grounded concepts, leading to embeddings that reflect a different notion of similarity. These distinctions explain why Flux, relying on T5 embeddings, exhibits different behavior from SD.
@@ Appendix C Attention Localization under Erasure Methods
Figure 9: Visualization of attention localization in concept erasure. The attention heatmaps confirm that existing erasure strategies [14, 20, 12, 3] share a common mechanism: attention localization. Localized regions corresponding to target concepts are suppressed, visually corroborating our theoretical analysis.
@@ Appendix D Why Naive Attention Maximization Diverges
$\begin{split}p_{i}=\frac{e^{z_{i}}}{\sum_{j=1}^{m}e^{z_{j}}},\quad i=1,\dots,m,\quad\\ \text{so }p\in(0,1)^{m},\ \sum_{i}p_{i}=1.\end{split}$ (12)
@@
$z_{k}-\max_{j\neq k}z_{j}\to+\infty\qquad\Longrightarrow\qquad p_{k}\to 1,\ p_{j\neq k}\to 0.$ (14)
@@
is structurally prone to divergence and vanishing gradients—precisely the recipe for the “attention divergence” we observe in the main text.
@@ Appendix E Flux Can Generate Consistent Content from Shuffled Prompts
The issue of overfitting arises from fixing the token_id of the target concept. To mitigate this, we introduce dynamic variation of the target token_id across training iterations. Prior work [14] speculated that randomly shuffling the prompt should not affect the generation quality of Flux. We extend this hypothesis with concrete evidence.
@@ Appendix F Baselines and Implementation Details
Erase AC [20]
ESD [12]
EraseAnything [14]
EAP [3]
ConceptPrune [6]
MACE [29]
Attack P4D [7]
UnlearnDiffAtk [57]
Ring-A-Bell [49]
Reason2Attack [54]
@@ F.1 Baselines for Attack Evaluation
We adopt UnlearnDiffAtk [57] as a representative white-box attack baseline, which leverages gradient-based noise prediction loss to optimize adversarial prompts, outperforming prior methods such as P4D [7]. For the black-box setting, we include Ring-A-Bell [49] and its enhanced variant Ring-A-Bell-Union, both of which construct a target concept direction from positive–negative prompt pairs and inject it into the prompt embedding to revive the forgotten concept, demonstrating stronger performance compared with approaches like QF-Attack [59]. For emerging methods, we consider the emerging LLM-based method Reason2Attack [54] as a representative of reasoning-driven attack strategies.
@@ F.2 Concept Erasure Methods for Evaluation
We select publicly accessible and reproducible concept erasure methods as victim models for evaluation. This includes classical approaches like ESD [12] and AC [20], as well as more recent methods, including EAP [3], ConceptPrune (CP) [6], MACE [29] and EraseAnything (EA) [14], the latter being the first approach tailored for rectified flow transformers. For ESD under nudity and violence settings, we fine-tune both non-cross-attention and cross-attention parameters with negative guidance factors of 1 and 3, respectively. We exclude UCE [13], as its overly aggressive removal severely distorts Flux outputs. All baselines and ablated models follow official implementations.
@@ H.2 Complete list of Entity, Abstraction, Relationship
This dataset is augmented on [14], covering more abstract and diverse test categories. The full list used in our experiments is presented in Table 8.
@@ H.3 Implementation Details of the Celebrity Benchmark
To construct a reliable benchmark for evaluating celebrity-related erasure and attack methods, we curate a refined subset from the CelebA dataset [27]. During this process, we deliberately exclude those individuals that Flux [dev] is unable to faithfully reconstruct. A manual inspection procedure is applied, where we compare synthesized outputs against their textual prompts and further supplement the pool with several well-known comic characters. This selection process ultimately yields a dataset of 100 celebrities, which we evenly divide into two groups: 50 designated for attack evaluation and 50 retained as control cases. The specific names of the celebrities used in our ablation study are listed in Table LABEL:table:_appendix_celebrity.
@@
For recognition tasks, we implement a lightweight yet effective classification network based on MobileNetV2 [17] pretrained on ImageNet [8]. On top of the original architecture, we append a GlobalAveragePooling2D layer followed by a fully connected Softmax layer. Training is performed with the Adam optimizer using a fixed learning rate of 1e-4, and categorical cross-entropy is adopted as the loss function. For training dataset, we first generate 50 images per celebrity (fixed prompt and random seeds), amounting to a total of 5,000 images. We then randomly re-sample the dataset and partitioned it into training (80%) and testing (20%) splits.
@@ H.4 Additional Results
Our attack method applies broadly across modern rectified-flow transformers. In the main paper, we report results on Flux.1 [dev]. Here we provide additional evaluations on other rectified-flow models, including Flux.1 [schnell] and SD 3. Because Flux.1 [schnell] shares the same architecture as Flux.1 [dev], our method transfers directly without modification. For SD3, we similarly fine-tune the text-conditioning parameters inside its MM-DiT that mediate interactions between text and image tokens.
Benchmarking Against State-of-the-Art (SOTA). Figure 12 shows the result of benchmarking ReFlux against representative SOTAs erasure attacks on the I2P dataset with Flux.1 [schnell] backbone model. For evaluation, we adopt NudeNet with a detection threshold of 0.6, which is commonly regarded as a reasonable boundary for identifying sensitive content. It should be noted, however, that surpassing this threshold does not necessarily imply the actual exposure of body organs, but rather indicates that the generated image triggers the detector’s nudity confidence. Among baselines, UnlearnDiffAtk exhibits a notable failure mode when applied to rectified flow models: it frequently collapses into low-quality or distorted generations. Nevertheless, due to the coarse sensitivity of the NudeNet classifier, such degraded outputs are still often flagged as “nude”, inflating its measured attack success. By contrast, our approach not only achieves higher quantitative scores under the same detector but also maintains high-fidelity and semantically consistent generations, providing a more accurate and reliable benchmark of erasure robustness.
@@
Figure 13: More visualization on artistic style (“Pablo Picasso”) attacks from the [6] dataset. While AC [20] erasure removes the distinctive Picasso style, our method successfully restores the erased artistic patterns across diverse works.
Attacking Artistic Style Concepts. Artistic styles are representative benchmarks of abstract concepts, widely evaluated in concept erasure and attack. Here, we present more visualization results. Figure 13 presents results for Picasso-style prompts. The top row shows baseline generations from SD 3, which faithfully capture Picasso’s unique color palette, geometric distortions, and expressive brushwork. After applying AC [20] erasure (middle row), these stylistic signatures vanish almost entirely, producing outputs that resemble conventional photographic or illustrative imagery rather than Picasso’s style. Our attack (bottom row) effectively restores the erased style, reintroducing signature characteristics such as fragmented spatial composition, bold outlines, and vibrant color schemes. Importantly, the restored images not only achieve a high attack success rate but also preserve layout and semantic content from the erased generations (e.g., maintaining the same subject matter and scene structure). These results demonstrate that our method can reactivate highly abstract and global concepts, highlighting the robustness and generality of our approach. This also reveals that defenses like AC achieve only surface-level erasure, leaving deep conceptual residues that can be readily reawakened—“erased, but not forgotten”.
@@ H.5 User Study and VLM-based Evaluation
Human User Study. To obtain reliable human judgments, we design questionnaires that sampled from 7 evaluation categories: nudity, violence, artistic style, entity, abstraction, relationship, and celebrity. For each category, we randomly selected 5 sets of comparison results in our main experiments, yielding one complete questionnaire. Sensitive content such as nudity or violence is masked with bars or blur to ensure participant safety. In total, we recruited 20 non-artist participants, each of whom completed on average 6 questionnaires.
VLM-based Evaluation. We leverage the latest GPT-5444https://openai.com/index/introducing-gpt-5/ VLM as an automated evaluator, chosen for its strong reasoning ability, nuanced understanding of semantics, and robust visual grounding. For each of the 7 categories, we select 10 representative comparison sets and present them to GPT-5. The evaluation criteria are aligned with those used in the human study and specified in Table 10, which we format as structured user prompts.
As shown in the main text and Figure 17, both our human user study and the GPT-5 VLM-based evaluation reveal a consistent trend. Our method outperforms all baseline approaches across the four metrics. Although Ring-A-Bell occasionally reports slightly higher values on irrelevant preservation, this is primarily because it fails to reactivate the target concept, producing outputs that remain almost indistinguishable from the erased baseline. In contrast, traditional PGD-based methods such as UnlearnDiffAtk and P4D often suffer from poor generation quality on Flux, exhibiting typical diffusion artifacts such as grid-like patterns, structural collapse, and mode instability, which result in substantially lower image quality scores. By effectively restoring the target concept while preserving prompt fidelity and visual coherence, our approach achieves both stronger semantic reactivation and more stable generative behavior, demonstrating clear advantages in robustness and reliability.